OpenStudy (anonymous):

A swimmer, capable of swimming at a speed of 1.0 m/s in still water (i.e., the swimmer can swim with a speed of 1.0 m/s relative to the water), starts to swim directly across a 3.0-km-wide river. However, the current is 0.91 m/s, and it carries the swimmer downstream. (a) How long does it take the swimmer to cross the river? (b) How far downstream will the swimmer be upon reaching the other side of the river?

OpenStudy (anonymous):

Part A) the swimmer still swims across at 1.0 m/s: the width of the river, x = 3 km = 3000 m The time taken by the swimmer to cross the river, t = x/v = 3000 m/1 = 3000 sec Part B) the current carries him downstream at a rate of .91 m/s: x = vt = 0.91m/s * 3000 s = 2730 m